### v1

Abstract
The title of a research paper communicates in a succinct style the main theme and, sometimes, the findings of the paper. Coming up with the right title is often an arduous task, and therefore, it would be beneficial to authors if title generation can be automated. In this paper, we fine-tune pre-trained and large language models to generate titles of papers from their abstracts. We also use ChatGPT in a zero-shot setting to generate paper titles. The performance of the models is measured with ROUGE, METEOR, MoverScore, BERTScore and SciBERTScore metrics. Our observations indicate that AI-generated paper titles are generally accurate and appropriate.
Keywords:
Natural language generation pre-trained language models large language models evaluation scholarly publishing
1 Introduction
One of the most important aspects in writing a research paper is coming up with a concise and informative title for the paper. In academic publishing, a title should be not only concise and informative but also engaging, ensuring that authors quickly grasp the paper’s insights. At a first glance, readers should gain a clear understanding of the contributions of the paper. Effective title contain keywords describing the reported research, and sentences in the paper containing words in the title typically reflect the paper’s main theme. Many studies have been devoted to exploring the connection between the title of a paper and its readership as captured by citations and downloads [1, 2, 3]. Short titles reportedly attract more citations [2].
Given the tremendous success of neural language models, especially large language models (LLMs), in various natural language processing (NLP) tasks, it is natural to ask if titles of scientific papers can also be generated by these models. Given the expectation that a title should capture the key import of a paper and yet be a short sequence of words that appeal to a large readership, the task is a challenging one. Title generation can be considered as a special case of abstractive text summarization that aims to distil out the most crucial information form a document.
In this paper, we aim to generate titles of research papers from their abstracts using deep neural models. In particular, we choose several pre-trained transformer models which we fine-tune with a dataset of abstracts and titles. We experiment with both the pre-trained and fine-tuned versions of LLaMA-3 LLM. We also use ChatGPT-3.5 in a zero-shot setting to generate titles from abstracts using prompts. We believe that machine generated titles could be very useful for non-native English speakers who find it difficult to quickly construct a suitable title for their papers, and also for novice researchers making their foray into the field. The generated titles can then be refined by them to more suit their own style and requirement.
The main contributions of this paper are:
  1. 1.
    We fine-tuned pre-trained models, namely, T5-base [4], BART-base [5], and PEGASUS-large [6] with CSPubSum dataset to automatically generate titles of research papers. We have also used the open large language model LLaMA-3-8B consist of 8 billion parameters [7][8]. We have contrasted its generation quality with and without fine-tuning it for this task. We have also also used ChatGPT 3.5 in a zero-shot setting to generate title from abstracts.
  2. 2.
    We have also curated a dataset of abstracts and titles, called LREC-COLING-2024, on which we evaluate our fine-tuned models without further training.
  3. 3.
    We evaluate the performance of the models using ROUGE [9], METEOR [10], MoverScore [11], BERTScore [12] and SciBERTScore metrics. We introduce SciBERTScore in this paper as a variant of BERTScore. We also evaluate the factual accuracy of the model-generated titles at the entity level by employing metrics such as precision-source and F1-target, which were introduced in a study by [13]. Manual evaluation of a small subset of the generated titles has also been carried out.
  4. 4.
    Our experiments show that PEGASUS-large outperforms the other models in terms of the chosen metrics. Notably, PEGASUS-large has fewer parameters compared to LLMs like GPT-3.5 and LLaMA.
  5. 5.
    A demo that allows the user to choose a pre-trained language model to generate a title from the abstract of a research paper is hosted at https://title-generation-researchpapers.onrender.com/. Additionally, on Hugging Face, we have publicly released the fine-tuned models [^1] https://huggingface.co/TRnlp as well as the LREC-COLING-2024 corpus [^2] https://huggingface.co/datasets/TRnlp/LREC-COLING-2024-Abstract-Title.
2 Literature survey
Automatic text summarization has a long history. As far back as 1958, Luhn et al. [14] pioneered an extractive summarization technique that selects sentences based on the frequency of significant words, excluding common words, to summarize technical papers and magazine articles. Lloret et al. [15] developed the COMPENDIUM system for generating biomedical research abstracts, utilizing both extractive and hybrid extractive-abstractive models to evaluate their effectiveness and suitability for summarization.
The subsequent achievements of sequence-to-sequence models [16], attention-based encoders with beam search and RNN [17], and pointer-generator networks with coverage mechanisms [18] have significant attention to the abstractive text summarization. All are mostly applied on news dataset. The next phase of research in summarization and natural language processing (NLP) in general was propelled by the transformer architecture [19]. It led to the development of various pre-trained language models such as T5 [20], BART [21], and PEGASUS [22] that are initially trained in self-supervised manner on broad, general-purpose text datasets. Their extensive pre-training helps them capture linguistic patterns and knowledge, which can then be fine-tuned for specific tasks across various domains.
In the special case of title generation for a body of text, researchers have proposed methods to generate titles of news articles from their content. Tan et al. [23] proposed a coarse-to-fine approach, which first identifies the important sentences with a with hierarchical attention with a encoder-decoder framework. They tested their model using New York Times (NYT) and DUC-2004 dataset. Our perspective in this paper acknowledges that a title essentially acts as a summary of the content within the research paper [24]. Mishra et al. [25] have proposed a method to generate titles for academic papers. First, they have generated a pool of candidate titles using a pre-trained model, then selected the most suitable one from them, and finally refined it to ensure semantic and syntactic accuracy, thereby enhancing the representativeness of titles for given texts. They have used arXiv [^3] https://tinyurl.com/y9pu6xyp, ACL [26], and ICMLA [27] dataset. Recently, Liu et al. [28] pre-trained a transformer encoder, OAG-BERT, on computer science research papers (including its metadata) and used it to produce paper titles, which were found to be quite acceptable to researchers although extensive evaluation of the generated titles using automatic evaluation measures has not been done. In contrast, we use a diverse set of encoder-decoder and decoder-only transformers in this paper to generate titles from abstracts. We also perform a detailed evaluation of the generated titles. Our work also aligns with the contemporary research [29] [30] that explores the potential of ChatGPT and other artificial intelligence tools in education, research, and scholarly publishing.
3 Datasets
We have used the CSPubSum dataset provided by Collins et al. [31], which contains URLs of 10147 computer science publications from ScienceDirect [^4] https://www.sciencedirect.com/. We crawled the dataset and organized each example as a pair of abstract and title. The dataset is split into 8120 examples for training, 1014 examples for validation, and 1013 examples for testing. We crawled an additional set of 1000 examples from accepted papers at LREC-COLING 2024555https://aclanthology.org/events/coling-2024/. This dataset, which we hereafter refer to as LREC-COLING-2024, comprises pairs of abstract, title for each paper.
We only train the selected deep neural language models on the train subset of CSPubSum but test them on the test subset of CSPubSum and the corpus of LREC-COLING-2024. While CSPubSum focuses on diverse topics of computer science, they primarily belong to the pre-transformer [19] era since the impact of transformers was produced after 2017 when CSPubSum was released. We wanted to verify how well the trained models perform on a different and more recent dataset on which it is not explicitly trained. So we curated LREC-COLING-2024. The average number of tokens in a title in CSPubSum is 12 while it is 11 for LREC-COLING-2024, although a few titles are longer and some as small as two tokens. A few titles have multiple sentences; a title with two sentences is “Comparative statics effects independent of the utility function. When do we act the same way under risk?”. However, around 82% papers in CSPubSum and 90% papers in LREC-COLING-2024 have no more than 15 tokens in their titles. So when fine-tuning or performing inference with the pre-trained models for title generation, we set the maximum title length to 20 tokens (recollect that due to sub-word generation, token count in a model generally exceeds the raw token count). Another reason for the imposed token limit is that shorter titles are generally more attractive to readers [2].
4 Methodology
In this section, we detail the fine-tuning process of various pre-trained language models (PLMs) and large language models (LLMs). We have chosen the following models:
  1. 1.
    T5-base [4]: It is an encoder-decoder model which is a slight variation of the original Transformer model [19]. Formulating every text processing problem, including translation, question answering, and classification, as a “text-to-text” transformation problem, the same model is trained to perform these diverse tasks. To pre-train the model, random text spans are corrupted/dropped and the model is trained to generate them. T5-base contains 220M parameters. We fine-tune the model with the train subset of CSPubSum.
  2. 2.
    BART-base [5]: A denoising autoencoder, it combines Bidirectional and Auto-Regressive Transformers exemplified by BERT [21] and GPT [32], respectively. To train BART, the input text is first corrupted with a noising function, and then the model reconstructs the original text. It is particularly suitable for text generation problems. BART-base is configured with 139M parameters. Like T5-base, it is also fine-tuned on CSPubSum train set.
  3. 3.
    PEGASUS-large [6]: It is a Transformer-based encoder-decoder model that is trained on large text collections with an objective function specifically focused on summarization. In particular, pre-training the model involves masking important sentences from an input document and generating them as one output sequence. PEGASUS-large contains 568M parameters. Like the previous two models, we fine-tune PEGASUS-large on training examples of CSPubSum.
  4. 4.
    ChatGPT-3.5: We use the instruction and chat-tuned version of the GPT-3 family model, identified as gpt-3.5-turbo-1106, commonly known as ChatGPT 3.5666https://chatgpt.com/. It is a decoder-only model that has been pre-trained on massive text corpora and then further trained with instruction fine-tuning and reinforcement learning with human feedback. It is a large language model with 175B parameters. We use the prompt-based in-context learning setup where we simply prompt the model to generate a title given the abstract.
  5. 5.
    LLaMA-3-8B: We use instruction-tuned LLaMA-3-8B model with fine-tuning and without fine-tuning. The LLaMA series of models [7] are decoder-only transformer-based large language models trained exclusively on publicly available datasets (in contrast to the ChatGPT and the GPT series). We fine-tune the LLaMA-3 model (with 8B parameters) on the train subset of CSPubSum.
Note that although the above pre-trained models and LLaMA-3 come in multiple sizes, we choose the smallest sizes that are freely available due to constraints on our computational resources. Although the paper that proposed the PEGASUS model [6] discusses a base and a large model, we could only find the checkpoints for PEGASUS-large in Hugging Face [^7] https://huggingface.co/google/pegasus-large and so, we consider this version. Although GPT-4 is more powerful than GPT-3.5, the former is only available to paid subscribers, and therefore, we do not use it for this research. However, we do expect that, like other deep learning applications, using larger models will enhance the quality of the generated titles.
4.1 Data processing
We removed only extra spaces from the documents in both datasets. We only retain examples where the abstract length is at least twenty tokens and the paper title length is at least three tokens. Since we impose a limit of 20 tokens on generated title length, we stipulate the abstract to be longer than 20 tokens so that the task can be treated as a text summarization problem.
4.2 Implementation details
We have chosen the following pre-trained models from the Hugging Face repository for fine-tuning on the CSPubSum dataset: T5-base [^8] https://huggingface.co/t5-base, BART-base [^9] https://huggingface.co/facebook/bart-base, PEGASUS-large [^10] https://huggingface.co/google/pegasus-large . Fine-tuning was performed for 5 epochs. We used batch size of 32, learning rate of 4e-5, and the metric_for_best_model as ROUGE1-F1.
We use the following prompt to generate titles using LLMs, namely ChatGPT-3.5 and LLaMA-3-8B,:
Create a concise title from this abstract using at most 20 tokens,
highlighting the main contributions and focus. <ABSTRACT>
Again during evaluation, both ChatGPT-3.5 and LLaMA-3-8B models were given the above prompt without any in-context examples. Unlike ChatGPT-3.5, LLaMA-3-8B can be trained. But fine-tuning LLaMA-3-8B is an extremely computationally expensive task. Therefore, we use the parameter-efficient fine-tuning technique called Low-Rank Adaptation (LoRA). In particular, we have taken the model from Hugging Face, LLaMA-3-8B  [^11] https://huggingface.co/unsloth/llama-3-8b-bnb-4bit, then loaded it in 4-bit precision to save memory, and finally, fine-tuned it for 5 epochs using a learning rate of 4e-5, train batch size of 32 and eval batch size of 1, rank of the adaptation matrices = r=16 , lora_alpha = 16 (Scaling Factor), and using peft_config for loading LoRA. For fine-tuning and evaluation after fine-tuning, we have used the same prompt that we mentioned above. For comparison, we also include results obtained with LLaMA-3-8B that is not fine-tuned on this dataset, i.e., it is used in zero-shot in-context learning without fine-tuning. We denote this variant as LLaMA-3-8B. In case of ChatGPT-3.5, we use a temperature setting of 0.3, to generate titles based on the abstracts.
For all models, we have set the maximum number of input tokens (i.e., the abstract length) to 512 and output tokens (i.e., the title length) to 20. In all cases, we have measured the memory and compute power consumption using the WandB tool [^12] https://wandb.ai/site. The fine-tuning of the LLaMA-3-8B model took 1 hour 43 minutes for 5 epochs. In the context of pre-trained language models, PEGASUS-large model took 23 minutes, BART-base model took 7 minutes, and T5-base took 13 minutes for 5 epochs of fine-tuning. Training and fine-tuning of the models were done on Tesla A100-SXM4-40GB Colab Pro+ that supports GPU-based training.
4.3 Evaluation metrics
We used commonly used automatic text summarization evaluation metrics, including ROUGE [9], METEOR [10], MoverScore [11], BERTScore [12] and SciBERTScore [33], to assess the quality of generated title with the author-written title. ROUGE scores measure $n$-gram overlap between the generated and ground-truth titles. We use ROUGE-1, ROUGE-2, and ROUGE-L where the first uses unigram overlap, the second bigram overlap and the last compares the longest common subsequence between the generated title and the golden title. METEOR measures sentence-level accuracy based on the alignment between the generated text and the reference text. In contrast, MoverScore, BERTScore, and SciBERTScore aim to measure the semantic similarity between the output and the true title, using latent representations (embeddings) of the texts. MoverScore combines Word Mover’s Distance [34] and contextualized embeddings of the output and the ground-truth titles for semantic matching [11]. BERTScore calculates the cosine similarity between the BERT embeddings of the words in the two text sequences [12]. Since the current application is in the domain of computer science, we also develop a variant of BERTScore where we use SciBERT [33] to generate the embeddings, then we compute their cosine similarity as in BERTScore; we call the modified metric, SciBERTScore.
These metrics are, however, inadequate to quantify factual consistency [35], which in our case connotes whether the entities present in the generated titles are precisely from the given abstract, and how much the entities in the generated title overlap with those in the author-written one. Hence, we have also used three new metrics, introduced by [13], to evaluate the factual consistency of the generated title.
Let us call the author-written title as the target $t$, the model-generated title as the hypothesis $h$ and the input abstract as the source $s$. We define $\mathcal{N}(t)$ as the count of named-entities in the target (author-written title) and $\mathcal{N}(h)$ as the count of named-entities in the hypothesis (model-generated title). $\mathcal{N}(h\cap s)$ is the number of entity matches across the generated title and the input source document (abstract). Some named-entities in the title span multiple words. If even some words of a named entity in the title occur in the abstract, we consider it as a successful entity match. We used the scispaCy model en_core_sci_sm [^13] https://allenai.github.io/scispacy/ to identify entities. Precision-source, defined as $prec_{s}=\mathcal{N}(h\cap s)/\mathcal{N}(h)$, is a metric that is used to determine the intensity of hallucination in relation to the input source document (abstract). Note that $prec_{s}$ represents the percentage of entities mentioned in the generated title that can be retrieved from the input source document (abstract). Low $prec_{s}$ indicates the possibility of hallucination in the generated title. However, $prec_{s}$ does not capture the generated title’s entity-level correctness in relation to the author-written title. Therefore, recollecting that the author-written title is the target $t$, we define target-level entity accuracy in terms of precision-target $prec_{t}=\mathcal{N}(h\cap t)/\mathcal{N}(h);$ recall-target $recall_{t}=\mathcal{N}(h\cap t)/\mathcal{N}(t);$ and F1-target $F1_{t}=\frac{2*(recall_{t}*prec_{t})}{recall_{t}+prec_{t}}.$ Here, $\mathcal{N}(h\cap t)$ represents the number of matched named-entities in the generated title and the author-written title. Note that we have calculated the above mentioned precision-source and precision-target in two ways. One is to consider the entity mentioned in each document (which may be the source $s$ or target $t$ or hypothesis $h$) as a set so that multiple occurrences of an entity in a document are equivalent to a single occurrence. The other is to consider the entity mentioned in a input source (abstract) as a list; here, if a metric is defined as $\mu=\mathcal{N}(x\cap y)/\mathcal{N}(x)$, then for each entity mention in $x$, we check if it occurs in $y$, and if so, increment the intersection count $\mathcal{N}(x\cap y)$ by unity; this approach is followed in [13]. In the first approach, we denote the metrics as $prec_{s}^{U}$, $prec_{t}^{U}$, $recall_{t}^{U}$, and $F1_{t}^{U}$ ($U$ indicates that only unique entity mentions are considered). In the second, we represent them as $prec_{s}^{NU}$, $prec_{t}^{NU}$, $recall_{t}^{NU}$, and $F1_{t}^{NU}$ ($NU$ denoting not unique).
5 Results
5.1 Quantitative comparison of different fine-tuned models
In this sub-section, we report the results of experiments on the test dataset of CSPubSum dataset as well as LREC-COLING-2024 dataset. Table 1 shows the performance in terms of ROUGE, METEOR, and semantic metrics like MoverScore, BERTScore and SciBERTScore, for the CSPubSum test dataset. Evaluation with entity-level factual consistency metrics precision-source($prec_{s}^{NU}$, $prec_{s}^{U}$), precision-target ($prec_{t}^{NU}$, $prec_{t}^{U}$), recall-target ($recall_{t}^{NU},recall_{t}^{U}$), and F1-target ($F1_{t}^{NU}$, $F1_{t}^{U}$) are shown in Table 2. The overall finding is that PEGASUS-large fine-tuned on the CSPubSum dataset achieves the highest scores for all the above metrics except precision-target metric. Table 3 and Table 4 show that the same PEGASUS-large model, that has been fine-tuned on the CSPubSum dataset, achieves the best performance on the LREC-COLING-2024 dataset except on BERTScore metric. Thus, the superlative performance of PEGASUS-large carries over to this dataset although it is not fine-tuned on it.
Table 1: Evaluation of all used models: F1-scores for ROUGE, METEOR are used for evaluating the both word-level matching and MoverScore, BERTScore, and SciBERTScore are used to compute the semantic fidelity of the generated titles for the CSPubSum dataset. All scores in percentage (%).
Model Name ROUGE-1 ROUGE-2 ROUGE-L METEOR MoverScore BERTScore SciBERTScore T5-base 44.25 25.04 38.92 38.36 38.09 89.9 76.06 BART-base 45.7 25.97 40.11 39.37 39.75 90.21 76.89 PEGASUS-large 46.75 27.13 40.67 42.61 40.43 90.35 76.93 LLaMA-3-8B 28.4 12.58 24.6 27.17 21.42 86.34 66.65 LLaMA-3-8B 40.8 21.23 36.57 34.5 37.02 89.99 76.41 ChatGPT-3.5 42.81 21.16 36.55 35.12 37.39 88.66 76.28
Table 2: Evaluation of all used models: precision-source, precision-target, recall-target and F1-target in terms of $prec_{s}^{NU}$, $prec_{s}^{U}$, $prec_{t}^{NU}$, $prec_{t}^{U}$, $recall_{t}^{NU}$, $recall_{t}^{U}$, $F1_{t}^{U}$ and $F1_{t}^{NU}$ scores are used for evaluating the factual consistency of the generated title for CSPubSum dataset. All scores in percentage (%).
Model Name $prec_{s}^{NU}$ $prec_{s}^{U}$ $prec_{t}^{NU}$ $recall_{t}^{NU}$ $F1_{t}^{NU}$ $prec_{t}^{U}$ $recall_{t}^{U}$ $F1_{t}^{U}$ T5-base 97.1 97.08 59.3 51.82 52.38 59.08 51.58 52.17 BART-base 97.44 97.39 61.52 52.84 53.91 61.35 52.56 53.72 PEGASUS-large 98.13 98.08 60.39 56.49 55.39 60.17 56.21 55.18 LLaMA-3-8B 67.7 67.7 39.19 40.48 36.81 38.96 39.98 36.48 LLaMA-3-8B 78.95 78.93 57.54 46.99 49.12 57.44 46.97 49.08 ChatGPT-3.5 92.73 92.73 59.79 50.95 51.7 59.79 51.01 51.74
Table 3: Evaluation of all used models: F1-scores for ROUGE, METEOR, MoverScore, BERTScore and SciBERTScore for LREC-COLING-2024 dataset. All scores in percentage (%).
Model Name ROUGE-1 ROUGE-2 ROUGE-L METEOR MoverScore BERTScore SciBERTScore T5-base 46.84 28.7 41.69 39.88 39.6 88.71 76.05 BART-base 46.87 27.66 41.89 38.93 39.61 88.84 75.94 PEGASUS-large 49.85 30.51 43.93 43.23 41.66 89.1 76.74 LLaMA-3-8B 32.92 16.66 27.66 30.61 25.68 86.77 67.42 LLaMA-3-8B 45.3 26.53 40.51 38.18 38.93 88.83 76.14 ChatGPT-3.5 45.16 23.97 38.88 37.45 38.85 89.54 75.64
Table 4: Evaluation of all used models: precision-source, precision-target, recall-target and F1-target in terms of $prec_{s}^{NU}$, $prec_{s}^{U}$, $prec_{t}^{NU}$, $prec_{t}^{U}$, $recall_{t}^{NU}$, $recall_{t}^{U}$, $F1_{t}^{U}$ and $F1_{t}^{NU}$ scores are used for evaluating the factual consistency of the generated title for LREC-COLING-2024 dataset. All scores in percentage (%).
Model Name $prec_{s}^{NU}$ $prec_{s}^{U}$ $prec_{t}^{NU}$ $recall_{t}^{NU}$ $F1_{t}^{NU}$ $prec_{t}^{U}$ $recall_{t}^{U}$ $F1_{t}^{U}$ T5-base 97.44 97.39 63.06 59.24 57.17 62.8 58.88 56.87 BART-base 96.24 96.23 64.3 59.21 57.78 64.11 58.92 57.56 PEGASUS-large 97.87 97.85 66.32 64.64 61.47 66.15 64.30 61.28 LLaMA-3-8B 74.6 74.52 45.29 48.16 43.48 45.02 47.64 43.13 LLaMA-3-8B 81.57 81.55 62.18 55.99 55.19 62.13 55.95 55.16 ChatGPT-3.5 91.98 91.98 60.87 57.66 55.56 60.84 57.69 55.56
We believe the reason PEGASUS-large wins the competition is that it is trained with an objective function that favors summarization, i.e., generating important lines that have been masked in the input document. Thus, our observations show that although the LLMs show excellent performance on a wide range of NLP tasks, smaller models fine-tuned on custom datasets may exhibit better performance on a given task.
A deep dive into the quantitative performance: Let us take a more nuanced view of the performance tables. Consider Table 1. It is immediately clear that the performance of LLaMA-3-8B (without fine-tuning) is considerably worse than that of the other models which are fine-tuned on the domain-specific data. This suggests that fine-tuning remains useful for attaining good model performance despite their extensive pre-training. In contrast, ChatGPT-3.5 performs better, even in a zero-shot setting, possibly because it is continuously used worldwide and as a result gets continually trained. A second observation is that although LLaMA-3-8B (fine-tuned) and ChatGPT-3.5 have smaller ROUGE and METEOR scores compared to the smaller pre-trained models, their performance is closer to that of the latter according to the semantic metrics like SciBERTScore; for example, SciBERTScore of all models except LLaMA-3-8B lie between 76.06 and 76.93. This indicates that all models (except LLaMA-3-8B) produce outputs with similar semantic similarity to the original titles. It also highlights the highly abstractive nature of the LLMs’ outputs, which, despite having low word overlap with the original titles, remain close in meaning to the original titles. Similar observations can be made from Table 3. Now let us analyze the entity overlap between the generated titles and the given titles and abstracts. Table 2 displays the figures for CSPubSum. First note that the precision-source as well as F1-target do not vary much whether or not we count multiple occurrences of an entity. Again, LLaMA-3-8B (without fine-tuning) shows the lowest scores for these metrics. The smaller pre-trained models perform similarly. Consider precision source $prec_{s}^{U}$: its large value indicates that the entities in the generated summary are mostly present in the input source which is the abstract here. $prec_{s}^{NU}=97.1$ for T5-base, $97.44$ for BART-base, and $98.13$ for PEGASUS-large. But $prec_{s}^{NU}$ is 92.73 for ChatGPT-3.5 and 78.95 for LLaMA-3-8B, showing that these models generate novel words more frequently. In contrast, the F1-target ($F1_{t}^{NU}$) lies in the 50’s for all models, meaning that there is only moderate overlap between the entities in the ground-truth title and the generated title. In particular, $F1_{t}^{NU}$ for LLMs is lower than that for the smaller models, again pointing to the abstractive nature of their output. We observe a similar pattern in Table 4 which reports the entity overlap statistics for the LREC-COLING-2024 dataset. Empirically, we did not observe noticeable hallucination in the generated titles but we intend to undertake a detailed study of this aspect in a future work.
5.2 Case studies
We show three representative examples of title generation one from CSPubSum (test subset) and two from LREC-COLING-2024 to illustrate the behavior of the models. Figure 1 shows the example from CSPubSum: the fine-tuned T5-base, BART-base, and PEGASUS-large models have generated titles that have significant similarity with the author-written titles. However, T5-base appears to generate a very long title. BART-base and PEGASUS-large generated the same titles. In contrast, LLaMA-3-8B without fine-tuning, produced a title that resembled a one-sentence summary and even that is incomplete; it is more of an extract from the abstract. After fine-tuning, the LLaMA-3-8B model generated an acceptable title. ChatGPT 3.5, using prompt-based techniques, successfully generated a title that also captures the essence of the paper and is similar to other generated titles and the author-written title. However, despite its stylistic flair, the evaluation metrics (in Table 1) indicate that the titles generated by ChatGPT-3.5 scored lower than those generated by the fine-tuned PEGASUS-large model. This discrepancy also highlights a limitation of the automated metrics in accurately evaluating titles generated in a highly abstractive and stylistic manner.
Author-written Title: “Comparative statics effects independent of the utility function. When do we act the same way under risk?”
fine-tuned T5-base: “Comparative statics effects independent of the utility function for portfolio choice and competitive firm under price uncertainty”
fine-tuned BART-base: “Comparative statics effects in the context of expected utility”
fine-tuned PEGASUS-large: “Comparative statics effects in the context of expected utility”
without fine-tuning LLaMA-3-8B: “The author proposes a methodological approach that enables comparative static analysis of various economic models, irrespective of the”
fine-tuned LLaMA-3-8B: “Comparative statics analysis: A new approach”
prompt based ChatGPT-3.5: “Comparative statics effects on expected utility in decision-making”
Figure 1: Input is an abstract from CSPubSum dataset. Title generated by each of the seven models is shown. Paper taken from https://www.sciencedirect.com/science/article/abs/pii/S037722171500586X
Now let us look at Fig. 2 which shows the outputs generated by the models for an example from LREC-COLING-2024 dataset. In case of BART-base, the generated title is incomplete. This is due to the hard limit on the output token count. We have observed this problem with other models, too. A simple workaround (without retraining the models) is to increase the output token limit during inference, which sometimes leads to a grammatically correct and complete title. While T5-base and fine-tuned LLaMA-3-8B generate a mostly correct title, they miss the “low-resource” setting which is correctly captured by PEGASUS-large. ChatGPT-3.5 generates almost the same title as that of PEGASUS-large except that it uses the abbreviation “NMT”, and abbreviations may not be desirable in a title.
Author-written Title: “A Reinforcement Learning Approach to Improve Low-Resource Machine Translation Leveraging Domain Monolingual Data”
fine-tuned T5-base: “Reinforcement Learning Domain Adaptation for Neural Machine Translation”
fine-tuned BART-base: “A novel Reinforcement Learning Domain Adaptation method for Neural Machine Translation in the low-”
fine-tuned PEGASUS-large: “Reinforcement learning domain adaptation for Neural Machine Translation in the low-resource domain”
without fine-tuning LLaMA-3-8B: “Reinforced Domain Adaptation Method for Low Resource Neural Machine Translation”
fine-tuned LLaMA-3-8B: “Reinforcement learning domain adaptation for neural machine translation”
prompt based ChatGPT-3.5: “Reinforcement Learning Domain Adaptation for Low-Resource NMT”
Figure 2: Input is an abstract from LREC-COLING-2024 dataset. Title generated by each of the seven models is shown. Paper taken from https://aclanthology.org/2024.lrec-main.132/
In the example in Fig. 3, PEGASUS-large and fine-tuned LLaMA-3-8B have generated exactly the same title as the author-written one. A quick reading of the abstract of the input paper indicates that T5-base has captured an important aspect of the proposed model in the paper, namely the “entity abstraction approach”; however, the model has other characteristics and singling out one might not be appropriate. The title generated by BART-base is logically incomplete as it misses the part “for entailment tree generation”. LLaMA-3-8B without fine-tuning struggles to generate a succinct title. But ChatGPT-3.5 does generate an interesting title which captures the very purpose of the proposed model: “improving AI explanations”
Author-written Title: “A Logical Pattern Memory Pre-trained Model for Entailment Tree Generation”
fine-tuned T5-base: “An entity abstraction approach for logical pattern memory pre-trained models”
fine-tuned BART-base: “Logical pattern memory pre-trained model”
fine-tuned PEGASUS-large: “A logical pattern memory pre-trained model for entailment tree generation”
without fine-tuning LLaMA-3-8B: “The proposed method addresses the limitations of previous approaches by incorporating an external memory structure to capture the latent representations”
fine-tuned LLaMA-3-8B: “Logical pattern memory pre-trained model for entailment tree generation”
prompt based ChatGPT-3.5: “Improving AI Explanations with Logical Pattern Memory Pre-trained Model”
Figure 3: Input is an abstract from LREC-COLING-2024 dataset. Title generated by each of the seven models is shown. Paper taken from https://aclanthology.org/2024.lrec-main.68/
5.3 Manual evaluation
We selected 20 papers, ten from each dataset (note that we use the test subset in case of CSPubSum) for human evaluation. A human annotator working in the field of NLP was asked to choose the most appropriate title among the generated ones. It is found that for CSPubSum dataset, in $80\%$ of the cases, the title generated by PEGASUS-large is most preferred, while for LREC-COLING-2024, in 50% cases, the output of PEGASUS-large wins while in 40% cases, the title generated by ChatGPT-3.5 is considered the best.
6 Demo
A demo of the application is hosted at https://title-generation-researchpapers.onrender.com/ where the user can input an abstract into a text box, select a suitable fine-tuned language model, and the maximum token count for the title, and obtain a title generated by the model. We have found that if the generated title is incomplete, increasing the token count (to say, 25 or 30) generally produces a correct and complete title.
Figure 4: Graphical user interface of our pre-trained language model-based title generation application.
7 Conclusion
We applied six different pre-trained/LLM models to generate titles of research papers. The fine-tuned PEGASUS-large model has achieved the best performance on the CSPubSum dataset as well the LREC-COLING-2024 dataset. Our extensive experiments suggest that pre-trained and large language models can generate engaging and suitable titles for research papers. The success of PEGASUS-large suggests that lighter pre-trained language models, fine-tuned on domain-specific datasets, may be satisfactory and even more suitable for this task than large language models with billions of parameters and substantial computational requirements. The high scores achieved by the models on LREC-COLING-2024 dataset on which they are not fine-tuned indicates that fine-tuning on a related scholarly dataset suffices to generate acceptable performance. Although the generated titles are generally of high quality, authors may further refine them to better suit their papers. This tool can be particularly useful for non-native English speakers and novice researchers. In future, we would like to analyze in more detail the differences between author-assigned and AI-generated titles and explore techniques to assess the quality of the abstractive nature of the outputs from LLMs.